OpenStudy (anonymous):

The variable t represents the temperature at 4 p.m. Which equation would best model the word problem below? The temperature at the summit of Mt. Mansfield dropped 17°F between 4 p.m. and 11 p.m. The temperature at 11 p.m. was –11°F. What was the temperature at 4 p.m.? A. 17 + 11 = t B. 11 – 17 = t C. t – 17 = –11 D. 17 – t = –11

OpenStudy (anonymous):

can you help me plz

OpenStudy (anonymous):

@Castiel

OpenStudy (josedavid):

c

OpenStudy (castiel):



OpenStudy (castiel):

Imagine you're out and it's 20F and 3 hours later it's 15F. How much did the temperature change? Well 20-5=15. It changed 5F. Or you could have written 20-t=15, you transfer all numbers to one side and all t's to other and you get t=20-15 which is t=5